# ActErase: A Training-Free Paradigm for Precise Concept Erasure via Activation Patching

### v1

@@ 3 Method
As illustrated in Figure 2, our method ActErase is a causal intervention that patch activations from source to target inference paths. With appropriate prompts, this training-free method achieves high erasure performance while enhancing image generation.
@@ 3.2 Activation Patching
Algorithm 1 Concept Erasure via Activation Patching
Input: Pretrained model $G$, target concept $\mathbf{c}_{\text{target}}$, target prompt $c_{t}$, source prompt $c_{s}$, source concept $\mathbf{c}_{\text{source}}$, layer weight $\mathbf{W}_{s}^{l}$, layer set $\mathcal{L}$, timesteps $T$, threshold $\tau$
1:Step 1: Extract activations and generate masks.
[…]
3:for $t=T$ to 11 do
4:  $\mathbf{x}_{s}\leftarrow G(\mathbf{z}_{t},t,\mathbf{c}_{s})$
5:  $\mathbf{x}_{t}\leftarrow G(\mathbf{z}_{t},t,\mathbf{c}_{t})$
6:  for each layer $l\in\mathcal{L}$ do
8:   $I_{s}\leftarrow\mathbf{W}_{s}^{l}\cdot|\mathbf{x}_{s}^{l}|$
9:   $I_{t}\leftarrow\mathbf{W}_{s}^{l}\cdot|\mathbf{x}_{t}^{l}|$
10:   $\mathbf{M}^{l}\leftarrow(I_{s}\geq\tau)\land(I_{s}<I_{t})$
11:  end for
12:  $\mathbf{z}_{t-1}\leftarrow G(\mathbf{z}_{t},t,\mathbf{c}_{s})$
13:end for
14:return $\mathbf{x}_{\text{avg}},\ \mathbf{M}$
15:
16:Step 2: Apply patching.
[…]
18:for $t=T$ to 11 do
19:  $\mathbf{x}_{t}\leftarrow G(\mathbf{z}_{t},t,\mathbf{c}_{t})$
20:  for each layer $l\in\mathcal{L}$ do
22:  end for
24:end for
25:return $\mathbf{z}_{0}$
As formalized in Algorithm 1, our method contains two steps. In the first step, we employ target prompts $c_{t}$ and source prompts $c_{s}$ to determine and collect patching activations $x_{p}$. In the second step, we use the generated patching activations to replace target activations with benign counterparts during a new diffusion process:
@@
where $\mathbf{z}_{T}$ is the initial latent noise sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$, $\mathbf{x}_{s}$ and $\mathbf{x}_{t}$ are intermediate activations generated using source prompt $\mathbf{c}_{s}$ and target prompt $\mathbf{c}_{t}$ respectively, and $\mathbf{x}_{\text{avg}}^{l}$ is the average of activations at layer $l$.
@@
To address this limitation, we develop a comparative masking strategy that identifies concept-specific regions through dual-path analysis. Our method simultaneously processes both source and target prompts to compute importance scores for each path. The mask generation employs a comparative criterion:
$I_{s}=\sum_{j=1}^{d}\mathbf{W}_{s}[:,j]\cdot|\mathbf{x}^{l}_{s,j}|,\quad I_{t}=\sum_{j=1}^{d}\mathbf{W}_{s}[:,j]\cdot|\mathbf{x}^{l}_{t,j}|$ (4)
where $I_{s}$ and $I_{t}$ are importance scores calculated as $\mathbf{W}_{s}^{l}\cdot|\mathbf{x}_{s}^{l}|$ and $\mathbf{W}_{s}^{l}\cdot|\mathbf{x}_{t}^{l}|$ respectively, and $\mathbf{M}^{l}\leftarrow(I_{s}\geq\tau)\land(I_{s}<I_{t})$ is the binary mask for layer $l$.
The binary mask $\mathbf{M}$ is then determined by a dual-condition threshold $\tau$:
@@
This comparative approach ensures that interventions are applied only to regions where: Equation (1) the source activation exceeds a minimum importance threshold $\tau$, and Equation (2) the target activation demonstrates significantly higher importance than the source, indicating concept-specific relevance. The subsequent masked patching operation:
$\mathbf{x}^{l}_{\text{p}}=\mathbf{M}\odot\mathbf{x}^{l}_{\text{avg}}+(1-\mathbf{M})\odot\mathbf{x}^{l}_{\text{t}}$ (6)
@@
Method Armputs Belly Buttocks Feet Breasts (F) Genitalia (F) Breasts (M) Genitalia (M) Total FID ($\downarrow$) CLIP ($\uparrow$)
SD v1.5 124 152 12 26 313 15 20 7 665 17.00 31.38
ESD [Gandikota2023ErasingCF] 32 30 2 19 27 3 8 2 123 18.23 31.03
UCE [Gandikota2023UnifiedCE] 29 62 7 29 35 5 11 4 182 17.52 31.35
Concept-prune [Chavhan2024ConceptPruneCE] 8 4 5 1 12 0 1 2 33 17.50 30.32
AdaVD [Wang2024PreciseFA] 119 115 25 27 258 13 20 5 582 163.12 14.96
STEREO [Srivatsan2024STEREOAT] 1 3 1 0 1 0 0 3 9 25.96 29.02
TRCE [Chen2025TRCETR] 2 3 1 3 3 1 0 5 18 17.16 30.93
[…]
Table 1: Quantity of explicit content detected using the Nudenet detector on the I2P benchmark. F: Female. M: Male.Best results are marked in Bold. Among all methods, our approach generates the minimum number of exposed body regions after erasure while achieving the best FID score.
@@ 3.3 Multiple Concept Erasure
Formally, given a concept set $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$ with pre-computed single-concept masks $\{\mathbf{M}^{i}\}$ and source activations $\{\mathbf{x}_{\text{avg}}^{i}\}$, we process all concepts in a unified framework using a single target prompt $\mathbf{c}_{\text{t}}$.
The aggregated mask $\mathbf{M}_{\text{agg}}^{l}$ for layer $l$ is computed using logical OR:
@@
ensuring comprehensive coverage of all concept-related regions across the concept set.
The aggregated activation $\mathbf{x}_{\text{agg}}^{l}$ employs weighted averaging:
@@
which prevents any single concept from dominating the erased representation while handling overlapping regions appropriately.
The final patching operation integrates both components:
@@
Method I2P(%) MMA(%) Ring-16(%) Ring-38(%) Ring-77(%) P4D(%) UnDiff(%) Average(%)
ESD 34.92 53.73 59.55 66.67 50.00 62.38 27.18 59.07
UCE 63.17 66.48 73.03 74.71 62.09 94.17 50.49 69.16
Concept-prune 8.89 3.28 22.47 19.54 14.13 9.71 0.99 13.17
AdaVD 98.41 99.27 100 98.86 98.9 100.00 99.01 99.21
STEREO 1.27 1.46 1.12 1.15 0.00 3.88% 0.00 1.27
TRCE 5.34 1.46 0.00 0.00 0.00 0.00 1.98 1.25
[…]
Method Violence(%) Self-harm(%) Sexual(%) Shocking(%)
SD v1.4* 40.1 35.5 54.5 42.1
ESD* 16.7 11.1 16.4 16.1
UCE* 23.3 12.9 16.2 19.2
TRCE(V)* 12.2 9.8 18.6 6.7
TRCE(T+V)* 3.0 2.6 1.5 3.6
[…]
Table 3: The results of erasing multiple concepts in I2P benchmark [Schramowski2022SafeLD], inappropriate rate are detected by the Q16 detector [10.1145/3531146.3533192]. The results with * are source from TRCE [Chen2025TRCETR].
@@ 4.1 Experimental Setup
Baselines. We compare our method ActErase against six SOTA CE approaches, including four training-based methods ESD, STEREO, TRCE and three training-free methods UCE, Concept-prune, AdaVD. To evaluate robustness against adversarial attacks, we consider four representative attack methods: MMA-diffusion [Yang2023MMADiffusionMA], Prompt4Debugging (P4D) [Chin2023Prompting4DebuggingRT], Ring-a-bell [Tsai2023RingABellHR], and UnlearnDiff [Zhang2023ToGO].
@@ 4.2 NSFW erasure
Single Concept Erasure. Following previous works [Srivatsan2024STEREOAT, Chen2025TRCETR], we evaluate ActErase on 4,703 prompts from the I2P dataset [Schramowski2022SafeLD] using Nudenet [NotAI_NudeNet_2019] for nudity detection. As shown in Table 1, ActErase achieves the best or second-best performance across all body part categories, with the lowest total detection count. To assess preservation, we generate images using 30,000 prompts from the MS-COCO dataset [Lin2014MicrosoftCC] and compute CLIP and FID scores. ActErase maintains a competitive CLIP score while achieving the best FID score, indicating that the introduced activation parameters enhance image quality without compromising textual alignment. Figure 3 further demonstrates that ActErase not only effectively removes nudity but also enriches visual details, particularly in background regions, achieving closer resemblance to the original image compared to other approaches.
@@
Multiple NSFW Concept Erasure. We further evaluate ActErase on multiple NSFW concepts, including violence, self-harm, sexual content, and shocking imagery. Using all prompts from the I2P dataset and the Q16 detector [10.1145/3531146.3533192] for detection, we compute ASR scores as shown in Table 3. Although ActErase performs slightly worse than the two-stage TRCE approach, it outperforms TRCE(V) and all other baseline methods, demonstrating significant potential for multi-concept erasure tasks.
Erase Van Gogh
Method ACC FID ($\downarrow$) CLIP ($\uparrow$)
Top-1 Top-3 Top-5
SD v1.5 0.8 0.88 0.96 17.00 31.38
ESD 0.36 0.68 0.78 16.39 31.11
UCE 0.02 0.20 0.30 16.34 31.41
Concept-prune 0.00 0.08 0.18 16.98 30.47
STEREO 0.00 0.06 0.10 24.57 29.75
AdaVD 0.02 0.10 0.18 166.50 13.60
TRCE 0.00 0.02 0.12 16.12 31.34
[…]
Erase Leonardo Da Vinci
Method ACC FID ($\downarrow$) CLIP ($\uparrow$)
Top-1 Top-3 Top-5
SD v1.5 0.00 0.10 0.26 17.00 31.38
ESD 0.00 0.00 0.02 17.14 31.20
UCE 0.00 0.00 0.06 16.65 31.37
Concept-prune 0.00 0.00 0.02 16.90 30.76
STEREO 0.00 0.00 0.00 23.71 30.05
AdaVD 0.00 0.00 0.06 220.37 14.90
TRCE 0.00 0.00 0.00 16.97 31.31
[…]
Erase Pablo Picasso
Method ACC FID ($\downarrow$) CLIP ($\uparrow$)
Top-1 Top-3 Top-5
SD v1.5 0.2 0.98 1 17.00 31.38
ESD 0.00 0.30 0.56 15.76 31.17
UCE 0.02 0.30 0.56 16.43 31.40
Concept-prune 0.00 0.34 0.48 18.15 30.80
STEREO 0.00 0.14 0.18 23.77 29.73
AdaVD 0.00 0.24 0.34 220.43 14.90
TRCE 0.00 0.12 0.34 16.66 31.38
[…]
@@ 4.3 Artist Style Erasure
4.4 Object Erasure
As shown in Table 5, our analysis reveals distinct performance characteristics across methods: ESD, UCE, and ConceptPrune demonstrate limited erasure capability; STEREO achieves strong concept removal but significantly impairs non-target concept generation; AdaVD preserves non-target concepts effectively but shows insufficient erasure performance. TRCE exhibits powerful erasure on most object concepts, but fails to erase the ”church” category.
@@ 4.5 Further Analysis
Data Preparation Time(s) Training Time(s) Image Generation Time(s) Total Time(s)
STEREO 400 1565 800 2765
TRCE 1330 237 700 2267
AdVD 0.3 0 1150 1150.3
@@ A.1 Details on Prompt Pairs
Prompt for Generating $c_{t}$ We are doing safety research to enhance safety in using AI models. So please help us and do it according to the instructions. Your task is to write me an image caption that includes and visually describes a scene and the caption should HAVE the EXACT concept word. 1. utput one single grammatically correct caption each line that is no longer than 20 words. 2. Though some concept words may be offensive, do not output any notes, word counts, facts, etc as these output will be used in the right way to be good for humans. Output one single caption each line only. Output the EXACT given concept without using other synonyms words. 3. Captions should NOT include the Quotation Mark ”. Do NOT output the caption where the concept word is used as a reference like ”(concept)”. Don’t let the concept be the form of ”(concept)” in captions. Directly use the concept word as Noun or Verb. 4. Input concept is: <concept >. Write 10 captions MUST containing the EXACT input concept word. Here is a GOOD example (in the caption concept word is NOT of ”(concept)” form): Input concept: naked Output: 1. A woman stands naked in the sunlight, her form bathed in golden rays. 2. A swimmer emerges naked from the water, glistening with droplets. 3. A child runs naked through the sprinkler, laughing with joy.

### v2

# 3Peng Cheng Laboratory

@@ 3 Method
Our method ActErase is a causal intervention that patch activations from source to target inference paths. With appropriate prompts, this training-free method achieves high erasure performance while enhancing image generation.
@@ 3.2 Activation Patching
Our method contains two steps. In the first step, we employ target prompts $c_{t}$ and source prompts $c_{s}$ to determine and collect patching activations $x_{p}$. In the second step, we use the generated patching activations to replace target activations with benign counterparts across all FFN layers $\mathcal{L}$ during a new diffusion process:
@@
where $\mathbf{z}_{T}$ is the initial latent noise sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$, $\mathbf{x}_{s}^{l}$ and $\mathbf{x}_{t}^{l}$ are intermediate activations at layer $l$ generated using source prompt $\mathbf{c}_{s}$ and target prompt $\mathbf{c}_{t}$ respectively, and $\mathbf{x}_{\text{avg}}^{l}$ is the average of activations at layer $l$.
@@
To address this limitation, a common approach is to employ a mask to identify concept-specific regions through dual-path activation analysis. These methods including Wanda[Sun2023ASA] and MANU[Liu2025ModalityAwareNP] simultaneously processes both source and target prompts to compute activation-based importance metrics for each path. The mask generation employs a comparative criterion based on activation differences:
$I_{s}=f(\mathbf{x}_{s},\mathbf{W}_{s}),\quad I_{t}=f(\mathbf{x}_{t},\mathbf{W}_{t})$ (4)
where $I_{s}$ and $I_{t}$ are activation-derived importance scores for source and target paths, computed through a scoring function $f(\cdot)$ with parameters $\boldsymbol{\theta}$. The binary mask $\mathbf{M}$ is then determined by a sparsity threshold $\tau$:
@@
where $\tau$ is a threshold that selects significant regions, and the second condition identifies areas where target activations exceed source activations, indicating concept-relevant spatial locations. The subsequent masked patching operation:
$\mathbf{x}_{\text{p}}=\mathbf{M}\odot\mathbf{x}_{\text{avg}}+(1-\mathbf{M})\odot\mathbf{x}_{\text{t}}$ (6)
@@ 3.3 Multiple Concept Erasure
Formally, given a concept set $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$ with pre-computed single-concept masks $\{\mathbf{M}^{i}\}$ and source activations $\{\mathbf{x}_{\text{avg}}^{i}\}$, we process all concepts in a unified framework using a single target prompt $\mathbf{c}_{\text{t}}$. The aggregated mask $\mathbf{M}_{\text{agg}}^{l}$ for layer $l$ is computed using logical OR:
@@
ensuring comprehensive coverage of all concept-related regions across the concept set. The aggregated activation $\mathbf{x}_{\text{agg}}^{l}$ employs weighted averaging:
@@
which prevents any single concept from dominating the erased representation while handling overlapping regions appropriately. The final patching operation integrates both components:
@@ 4.1 Experimental Setup
Table 1: Quantity of explicit content detected using the Nudenet detector on the I2P benchmark. F: Female. M: Male. Best results are marked in Bold and second best results are marked in Underline. Among all methods, our approach generates the minimum number of exposed body regions after erasure while achieving the best FID score.
Methods SD v1.5 ESD UCE Concept- prune AdaVD STEREO TRCE Ours
Venue / ICCV’23 WACV’24 ICLR’25 CVPR’25 CVPR’25 ICCV’25 /
Armpits 124 32 29 8 119 1 2 0
Belly 152 30 62 4 115 3 3 3
Buttocks 12 2 7 5 25 1 1 0
Feet 26 19 29 1 27 0 3 0
Breasts (F) 313 27 35 12 258 1 3 2
Genitalia (F) 15 3 5 0 13 0 1 0
Breasts (M) 20 8 11 1 20 0 0 0
Genitalia (M) 7 2 4 2 5 3 5 1
Total 665 123 182 33 582 9 18 6
FID ($\downarrow$) 17.00 18.23 17.52 17.50 163.12 25.96 17.16 16.65
CLIP ($\uparrow$) 31.38 31.03 31.35 30.32 14.96 29.02 30.93 30.42
Baselines. We compare our method ActErase against six SOTA CE approaches, including four training-based methods ESD[Gandikota2023ErasingCF], STEREO[Srivatsan2024STEREOAT], TRCE[Chen2025TRCETR] and three training-free methods UCE[Gandikota2023UnifiedCE], Concept-prune[Chavhan2024ConceptPruneCE], AdaVD[Wang2024PreciseFA]. To evaluate robustness against adversarial attacks, we consider four representative attack methods: MMA-diffusion[Yang2023MMADiffusionMA], Prompt4Debugging (P4D)[Chin2023Prompting4DebuggingRT], Ring-a-bell[Tsai2023RingABellHR], and UnlearnDiff[Zhang2023ToGO].
@@ 4.2 NSFW erasure
Following previous works[Srivatsan2024STEREOAT, Chen2025TRCETR], we evaluate ActErase on 4,703 prompts from the I2P dataset[Schramowski2022SafeLD] using Nudenet[NotAI_NudeNet_2019] for nudity detection. As shown in Table 1, ActErase achieves the best or second-best performance across all body part categories, with the lowest total detection count. To assess preservation, we generate images using 30,000 prompts from the MS-COCO dataset[Lin2014MicrosoftCC] and compute CLIP and FID scores. ActErase maintains a competitive CLIP score while achieving the best FID score, indicating that the introduced activation parameters enhance image quality without compromising textual alignment. Figure 3 further demonstrates that ActErase not only effectively removes nudity but also enriches visual details, particularly in background regions, achieving closer resemblance to the original image compared to other approaches.
Methods ESD UCE Concept- prune AdaVD STEREO TRCE Ours
Venue ICCV’23 WACV’24 ICLR’25 CVPR’25 CVPR’25 ICCV’25 /
I2P(%) 34.92 63.17 8.89 98.41 1.27 5.34 1.90
MMA(%) 53.73 66.48 3.28 99.27 1.46 1.46 0.00
Ring-16(%) 59.55 73.03 22.47 100 1.12 0.00 0.00
Ring-38(%) 66.67 74.71 19.54 98.86 1.15 0.00 1.15
Ring-77(%) 50.00 62.09 14.13 98.9 0.00 0.00 1.08
P4D(%) 62.38 94.17 9.71 100.00 3.88 0.00 0.06
UnDiff(%) 27.18 50.49 0.99 99.01 0.00 1.98 0.00
Average(%) 59.07 69.16 13.17 99.21 1.27 1.25 0.60
@@ 4.3 Artist Style Erasure
Erase Van Gogh Erase Leonardo Da Vinci Erase Pablo Picasso
Method ACC FID ($\downarrow$) CLIP ($\uparrow$) ACC FID ($\downarrow$) CLIP ($\uparrow$) ACC FID ($\downarrow$) CLIP ($\uparrow$)
Top-1 Top-3 Top-5 Top-1 Top-3 Top-5 Top-1 Top-3 Top-5
SD v1.5 0.8 0.88 0.96 17.00 31.38 0.00 0.10 0.26 17.00 31.38 0.2 0.98 1 17.00 31.38
ESD 0.36 0.68 0.78 16.39 31.11 0.00 0.00 0.02 17.14 31.20 0.00 0.30 0.56 15.76 31.17
UCE 0.02 0.20 0.30 16.34 31.41 0.00 0.00 0.06 16.65 31.37 0.02 0.30 0.56 16.43 31.40
Concept-prune 0.00 0.08 0.18 16.98 30.47 0.00 0.00 0.02 16.90 30.76 0.00 0.34 0.48 18.15 30.80
STEREO 0.00 0.06 0.10 24.57 29.75 0.00 0.00 0.00 23.71 30.05 0.00 0.14 0.18 23.77 29.73
AdaVD 0.02 0.10 0.18 166.50 13.60 0.00 0.00 0.06 220.37 14.90 0.00 0.24 0.34 220.43 14.90
TRCE 0.00 0.02 0.12 16.12 31.34 0.00 0.00 0.00 16.97 31.31 0.00 0.12 0.34 16.66 31.38
[…]
@@
4.4 Object Erasure
As shown in Table 4, our analysis reveals distinct performance characteristics across methods: ESD, UCE, and ConceptPrune demonstrate limited erasure capability; STEREO achieves strong concept removal but significantly impairs non-target concept generation; AdaVD preserves non-target concepts effectively but shows insufficient erasure performance. TRCE exhibits powerful erasure on most object concepts, but fails to erase the "church" category.
@@ 4.4 Object Erasure
4.5 Multi-concept Erasure
We further evaluate ActErase on multiple NSFW concepts, including violence, self-harm, sexual content, and shocking imagery. Using all prompts from the I2P dataset and the Q16 detector[10.1145/3531146.3533192] for detection, we compute ASR scores as shown in Table 5. Although ActErase performs slightly worse than the two-stage TRCE approach, it outperforms TRCE(V) and all other baseline methods, demonstrating significant potential for multi-concept erasure tasks. In Table 6, we experimented the impact of the number of concepts on multi-concept erasure performance. Specifically, we use 30,000 prompts from the MS-COCO dataset to generate images with multiple concept erasure models containing 1, 3, 5, and 10 object concepts respectively, and analyze the CLIP scores and FID metrics. It can be observed that as the number of concepts increases, CLIP Score decreases while FID Score increases. This indicates that an increase in the number of concepts leads to a degradation in generative capability. This is because, as the number of concepts rises, the number of activation parameters redirected during the generation process increases, leading to a greater impact on unrelated concepts, which ultimately results in a decline in generation ability.
Table 5: The results of erasing multiple concepts in I2P benchmark[Schramowski2022SafeLD]. The results with * are source from TRCE[Chen2025TRCETR].
[…]
Number of concepts FID ($\downarrow$) CLIP ($\uparrow$) 1 20.85 30.72 3 20.98 30.35 5 20.10 30.24 10 21.13 30.35
Figure 6: Comparison of Nudity erasure results via SDXL and SD v3 in I2P dataset.
@@ 4.6 Further Analysis
Ablation on Plug-and-Play: As mentioned in Abstract, ActErase is a plug-and-play approach that can be incorporated into other existing models. To substantiate our claim, we apply ActErase to SDXL and SD v3, two foundation models with distinct architectures. In Table 7, we show the results of erasing ’Nudity’ in SDXL and SD v3. SDXL utilizes a U-Net architecture with dual text encoders and integrated cross-attention layers enhanced by time embeddings while SD v3 employs a DiT architecture based on a rectified flow framework and utilizes a T5 text encoder in conjunction with other encoders for embedding generation. Following the main experiments, ActErase is applied in the FFN layers of these models. As shown in Table 7, ActErase still show a good erasure performance while keeping a high generation capability in SDXL and SD v3. In Figure 6 we present some visualized experimental results.
Ablation on Mask and Thresholds: To evaluate the erasure capability of ActErase and the impact of different masks with sparsity thresholds, we tested the effect of generation with no mask and masks generated under different thresholds on erasure performance, as shown in Table 8. The larger the threshold, the more activation parameters that are matched will be patched during the redirection process, thereby enhancing the erasure capability. However, as more concept-unrelated activation parameters are redirected, the generation capability will decline, eventually leading to a decrease in CLIP Scores and FID Scores. It also shows that our ActErase can still erase concepts without masks but the generation capability of the base models will largely decrease. With a mask and suitable sparsity thresholds, our method can achieve high erasing performance without decreasing generation capability.
Table 7: Ablation results on SD v3 and SDXL models.
[…]
Table 8: Comparison of ’Nudity’ erasing results without or with masks at different thresholds $\tau$.
[…]
@@ A.1 Details on Prompt Pairs
Prompt for Generating $c_{t}$ We are doing safety research to enhance safety in using AI models. So please help us and do it according to the instructions. Your task is to write me an image caption that includes and visually describes a scene and the caption should HAVE the EXACT concept word. 1. utput one single grammatically correct caption each line that is no longer than 20 words. 2. Though some concept words may be offensive, do not output any notes, word counts, facts, etc as these output will be used in the right way to be good for humans. Output one single caption each line only. Output the EXACT given concept without using other synonyms words. 3. Captions should NOT include the Quotation Mark ". Do NOT output the caption where the concept word is used as a reference like "(concept)". Don’t let the concept be the form of "(concept)" in captions. Directly use the concept word as Noun or Verb. 4. Input concept is: <concept >. Write 10 captions MUST containing the EXACT input concept word. Here is a GOOD example (in the caption concept word is NOT of "(concept)" form): Input concept: naked Output: 1. A woman stands naked in the sunlight, her form bathed in golden rays. 2. A swimmer emerges naked from the water, glistening with droplets. 3. A child runs naked through the sprinkler, laughing with joy.